# GAPO: Group Adaptive Policy Optimization for Real-World Code Edit

### v1

@@ Introduction
With the rapid advancement of large language models (LLMs), artificial intelligence (AI)–assisted coding has emerged as a prominent subfield and practical application (Chen et al. 2021; Li et al. 2022), demonstrating proven value in improving software engineering efficiency (Peng et al. 2023; Yetiştiren et al. 2023) and serving as a tool-generation assistant for white-collar workers across diverse domains (Schick et al. 2023; Achiam et al. 2023). Most code LLMs undergo a post-training stage, and reinforcement learning (RL) — in particular, RL from learned or human preference–based rewards— is a widely used post-training method (Christiano et al. 2017; Wang et al. 2024).
Among RL methods, Group Relative Policy Optimization (GRPO) (Shao et al. 2024) and its variants are popular choices, known for their robustness across diverse scenarios (Liu et al. 2025c). The core characteristic of the GRPO family lies in its group-relative advantage computation, which removes the need for a critic model (or value function) by sampling a group of rollouts for each input prompt and calculating advantage values as normalized rewards relative to the mean reward within each group (Shao et al. 2024), as illustrated in Figure 1. However, in real-world code-editing scenarios with complex contexts, inter-module invocation relationships, and diverse user intents, input prompts can inevitably induce rollouts that contain outliers — rollouts that are both unpredictable (Frauenknecht et al. 2025; Wu et al. 2022) and noisy (Liu et al. 2025a). In such practical cases, the expected normal or symmetric reward distributions often shift toward left- or right-skewed forms (Moore, McCabe, and Craig 2009), a phenomenon frequently observed in practice (see Figure 7 for details). Given a normally distributed reward within the range $[0,1]$, when most rewards are greater than 0.5 (i.e., the mean exceeds 0.5), the practical reward distribution becomes left-skewed, as unpredictable outliers may appear anywhere within the range, but their impact is more pronounced in the lower-reward region (below 0.5). A similar pattern holds for right-skewed cases.
@@
To mitigate the negative impact of outliers, we propose Group Adaptive Policy Optimization (GAPO), a group-adaptive advantage computation method designed to enhance existing GRPO and its variants (e.g., DAPO (Yu et al. 2025)), as illustrated in Figure 1. In the original formulation, the mean treats all rewards uniformly—including outliers—which can distort the advantage calculation. To address this, we first adaptively identify an outlier-free sub-region within all rewards by reformulating the problem as the classical highest-density interval (HDI) detection task from statistics (O’Neill 2022), solved using an adapted sliding-window scan algorithm. Since reward values are concentrated in the highest-density region, which is less affected by outliers, this enables us to determine an adaptive outlier-free sub-region for each input prompt. Subsequently, to further enhance robustness against outliers, we select the median within this region as the adaptive $Q$, replacing the original mean in the computation of group-relative advantage. As shown in Figure 1, the adaptive $Q$ provides additional benefits. In left-skewed distributions, GAPO generates more negative rollouts than GRPO/DAPO, since $Q$ exceeds the mean and most rewards fall below it, producing negative advantages. This is reasonable, as left-skewed distributions usually correspond to relatively easy problems with high rewards, where negative samples improve generalization (Mu et al. 2025; Zhu et al. 2025). Conversely, for right-skewed distributions (relatively hard problems), GAPO promotes more specialized learning, enhancing accuracy on these challenging cases.
@@
where $\hat{e}$ and $e^{*}$ represent two code snippets (LLM output and ground truth, respectively), $ed(\hat{e},e^{*})$ is computed using the classical dynamic programming algorithm (Lcvenshtcin 1966), and $l(\cdot)$ returns the length of a code snippet.
@@
GRPO is one of the most widely adopted methods in both industry and research within the field of reinforcement learning (Pennino et al. 2025; Li et al. 2025a). It is particularly effective for addressing the optimization problem in Eq. 1. A key advantage of GRPO is that it eliminates the need for a separate value function approximation; instead, it calculates advantages using the average reward across multiple sampled responses (rollouts) generated from the same input prompt. When using GRPO from an RL view, we can rewrite Eq. 1 to be
@@
where $\pi_{\theta}$, $\pi_{\theta_{old}}$, and $\pi_{ref}$ are the updating policy model, old policy model, and reference model, respectively. For each prompt $q$, GRPO samples a group of $G$ edits $\{e_{1},e_{2},\ldots,e_{G}\}$ from the old policy $\pi_{\theta_{old}}$. Here, $\epsilon$ and $\beta$ are hyperparameters that control the optimization behavior, and $\mathbb{D}_{\mathrm{KL}}$ denotes the KL divergence between the current training policy and a fixed reference policy (see (Shao et al. 2024) for details).
@@
As demonstrated in our experiments (see Figure 7 for details), a substantial fraction of input prompts in real-world code editing scenarios yield skewed reward distributions. This skew arises from the inherent noise, complexity, and dynamism of practical environments—often manifested through outliers—which are unavoidable and unpredictable due to factors such as rich contextual dependencies, historical user behaviors, evolving intents, and other confounding variables. These outliers introduce noise into RL training (Frauenknecht et al. 2025; Hollenstein et al. 2022) when using Eq. 4, since Eq. 4 employs an aggregation function that uniformly considers all reward values. Common approaches to mitigating the influence of outliers include using quantile-based statistics (e.g., Q1, median, Q3), which are more robust than the mean in the presence of outliers (John 2015; Rousseeuw and Hubert 2011). However, as shown in Figure 7 and Figure 8, real reward distributions across different input prompts vary—being left-skewed, normal, or right-skewed—so a single quantile or mean cannot adapt effectively to diverse cases. To address this issue, we propose a group-adaptive advantage calculation approach to enhance existing GRPO and its variants (such as DAPO (Yu et al. 2025)); this novel RL method is termed Group Adaptive Policy Optimization (GAPO). The core idea is to identify the highest-density region of rewards, where most reward values concentrate, and the influence of outliers is minimal.
@@
Our GAPO method does not alter the objective of existing group-relative RL approaches; instead, it only modifies the advantage computation in Eq. 4. This design makes it simple to implement and plug-and-play with any RL framework, such as verl (Sheng et al. 2025), requiring only a few lines of code. Specifically, we redefine the advantage as
@@
To identify an outlier-free region that adapts to different reward distributions, we seek a sub-region with the highest density, which represents the majority of the data while ignoring outliers. This naturally leads us to a classical statistical problem: finding the highest-density interval (HDI) for a given probability mass (O’Neill 2022), which also identifies the narrowest interval. Formally, for a set of reward values, this can be expressed as
@@
For comprehensive evaluation, we consider nine large language models (LLMs) covering a diverse spectrum: Mistral-v0.3 (Jiang et al. 2023) (general-purpose, non-reasoning; 7B), Qwen2.5 (Qwen Team 2024) (general-purpose, non-reasoning; 3B, 7B), Qwen3 (Yang et al. 2025) (general-purpose, reasoning; 4B, 8B), Qwen2.5-Coder (Hui et al. 2024) (code-specialized, non-reasoning; 3B, 7B, 14B), and DeepSeek-Coder (Guo et al. 2024a) (code-specialized, non-reasoning; 6.7B). All these models are instruction-tuned versions downloaded from Hugging Face (Hugging Face 2025).
@@
Group Relative Policy Optimization (GRPO) (Shao et al. 2024) has emerged as one of the most influential RL algorithms for LLM post-training (Kumar et al. 2025; Patil and Jadon 2025; Lai et al. 2025), widely adopted in both academia and industry. Recently, Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al. 2025) has gained attention as a promising extension and popular successor to GRPO. Our proposed GAPO method focuses on enhancing the group advantage computation—the core mechanism of GRPO. Accordingly, we adopt both the classical GRPO and the newly introduced DAPO as our primary baselines. We demonstrate here that the adaptive advantage computation, at the heart of GAPO, is a generalizable framework that can be seamlessly integrated into both GRPO and DAPO, leading to consistent performance improvements.
@@
We use the popular verl framework (Sheng et al. 2025) for RL training, which is widely adopted in industry due to its scalability. We follow most of the default settings for GRPO and DAPO in verl with adaptive modifications for our code edit tasks, such as an input prompt length of 4096, an output length of 1024, a rollout batch size of 512, a training batch size of 32, 8 rollouts per iteration, and a total of 10 epochs. Our GAPO method is straightforward to implement, requiring only a few lines of code by modifying the compute_grpo_outcome_advantage function in the GRPO and DAPO implementations (details are provided in our released code), demonstrating its high compatibility. We use the exact match ($em$) as the evaluation metric. All results are reported as the average over three trials.
@@
We investigate the reward distributions of rollouts for all input prompts in the training set, with the resulting statistics shown in Figure 7. The pie chart indicates that most input prompts (68.4%) induce rollouts with approximately normal distributions, while a substantial portion (19.9%) exhibits skewed distributions, roughly evenly split between left-skewed and right-skewed. This skew arises from the inherent noise, complexity, and dynamism of real-world environments—often manifesting as unavoidable and unpredictable outliers driven by complex context, user history, evolving intents, and other confounding factors. When computing advantage values as in Eq. 4 for these skewed distributions, outliers can have a negative impact, introducing noise (Frauenknecht et al. 2025; Hollenstein et al. 2022). In contrast, our GAPO method employs an adaptive $Q$ value to locate the center of the dense region in the reward distribution, thereby mitigating the influence of outliers and reducing noise in the computation of advantage values.
@@
Furthermore, we visualize representative examples sampled from input prompts that exhibit four types of reward distributions in Figure 8, providing a closer look at the skewed cases. Noticeable gaps exist between the mean and median for the left-skewed, right-skewed, and approximately symmetric distributions, while outliers primarily appear in the left- and right-skewed cases. By further analyzing the relationship between the mean–median difference and the sign of the advantages computed in Eq. 4 and Eq. 5, we find that our GAPO method generates more negative rollouts than the original GRPO/DAPO in left-skewed distributions. This occurs because $Q$ is larger than the mean in such cases, causing most rewards to fall below $Q$ and thus produce negative advantages. This behavior is reasonable, as left-skewed distributions typically correspond to relatively easy problems with generally higher rewards, where more negative samples promote better generalization during training (Mu et al. 2025; Zhu et al. 2025). Conversely, for right-skewed distributions (relatively hard problems), GAPO encourages more specialized learning, leading to improved accuracy on these challenging cases. These trends align with our post-training goals.
@@
To address the instability and limited exploration of GRPO, recent studies have proposed improvements from multiple perspectives. For instance, (Wei et al. 2025) highlights GRPO’s gradient instability and mitigates it by dithering discrete reward signals with random noise. GMPO (Zhao et al. 2025) replaces the arithmetic mean with a geometric mean to reduce outlier sensitivity. Dr. GRPO (Liu et al. 2025b) corrects training bias by jointly considering response length and question difficulty. GSPO (Zheng et al. 2025) mitigates high variance in Mixture-of-Experts models via sequence-level importance sampling. Other works (Cui et al. 2025; Hao et al. 2025b) stabilize training and alleviate entropy collapse by regulating entropy dynamics.
Beyond token-level optimization, another line of work enhances GRPO stability through refined advantage computation. OPO (Hao et al. 2025a) achieves stable policy optimization via a variance-minimizing bias term. BNPO (Xia et al. 2025) adaptively normalizes rewards using a Beta distribution for low-variance gradient estimation. (Wang et al. 2025b) employs Kalman filtering to estimate latent reward baselines and uncertainty. GRPO-MA (Wang et al. 2025a) reduces variance by averaging multi-answer rewards, while (Wu et al. 2025a) introduces a group-wise $K$-quantile baseline for entropy-safe reasoning.
@@
Code editing (Li et al. 2023; Guo et al. 2024b; Nam et al. 2025) is more challenging than basic code refinement, as it depends not only on the user prompt but also on the coding context, cursor position, and historical edits. To tackle this challenge, Self-Edit (Zhang et al. 2023) fine-tunes a fault-aware neural editor in a generate-and-edit manner, improving code quality and accuracy. EDITLORD (Li et al. 2025b), in contrast, avoids direct prompting or fine-tuning; it first uses an LLM to extract transformation rules from training data and then trains a second LLM to apply these rules to code pairs. IterPref (Wu et al. 2025b) leverages offline preference learning for iterative debugging, enabling context-aware code editing based on user feedback. LEDEX (Jiang et al. 2024) automatically collects refinement trajectories and enhances LLMs’ self-debugging ability via supervised fine-tuning and reinforcement learning. Finally, RLEF (Gehring et al. 2024) performs code editing with real-time execution feedback, optimizing refinement through end-to-end reinforcement learning.

### v2

# GAPO: Robust Advantage Estimation for Real-World Code LLMs

@@ Introduction
With the rapid advancement of large language models (LLMs), artificial intelligence (AI)–assisted coding has emerged as a prominent subfield and practical application (chen2021evaluating; li2022competition), demonstrating proven value in improving software engineering efficiency (peng2023impact; yeticstiren2023evaluating) and serving as a tool-generation assistant for white-collar workers across diverse domains (schick2023toolformer; achiam2023gpt). Most code LLMs undergo a post-training stage, and reinforcement learning (RL) — in particular, RL from learned or human preference–based rewards— is a widely used post-training method (christiano2017deep; wang2024enhancing).
Among RL methods, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) and its variants are popular choices, known for their robustness across diverse scenarios (liu2025part). The core characteristic of the GRPO family lies in its group-relative advantage computation, which removes the need for a critic model (or value function) by sampling a group of rollouts for each input prompt and calculating advantage values as normalized rewards relative to the mean reward within each group (shao2024deepseekmath), as illustrated in Figure 1. However, in real-world code-editing scenarios with complex contexts, inter-module invocation relationships, and diverse user intents, input prompts can inevitably induce rollouts that contain outliers — rollouts that are both unpredictable (frauenknechtrollouts; wu2022plan) and noisy (liu2025noisyrollout). In such practical cases, the expected normal or symmetric reward distributions often shift toward left- or right-skewed forms (moore2009introduction), a phenomenon frequently observed in practice (see Figure 7 for details). Given a normally distributed reward within the range $[0,1]$, when most rewards are greater than 0.5 (i.e., the mean exceeds 0.5), the practical reward distribution becomes left-skewed, as unpredictable outliers may appear anywhere within the range, but their impact is more pronounced in the lower-reward region (below 0.5). A similar pattern holds for right-skewed cases.
@@
To mitigate the negative impact of outliers, we propose Group Adaptive Policy Optimization (GAPO), a group-adaptive advantage computation method designed to enhance existing GRPO and its variants (e.g., DAPO (yu2025dapo)), as illustrated in Figure 1. In the original formulation, the mean treats all rewards uniformly—including outliers—which can distort the advantage calculation. To address this, we first adaptively identify an outlier-free sub-region within all rewards by reformulating the problem as the classical highest-density interval (HDI) detection task from statistics (o2022smallest), solved using an adapted sliding-window scan algorithm. Since reward values are concentrated in the highest-density region, which is less affected by outliers, this enables us to determine an adaptive outlier-free sub-region for each input prompt. Subsequently, to further enhance robustness against outliers, we select the median within this region as the adaptive $Q$, replacing the original mean in the computation of group-relative advantage. As shown in Figure 1, the adaptive $Q$ provides additional benefits. In left-skewed distributions, GAPO generates more negative rollouts than GRPO/DAPO, since $Q$ exceeds the mean and most rewards fall below it, producing negative advantages. This is reasonable, as left-skewed distributions usually correspond to relatively easy problems with high rewards, where negative samples improve generalization (mu2025dissecting; zhu2025surprising). Conversely, for right-skewed distributions (relatively hard problems), GAPO promotes more specialized learning, enhancing accuracy on these challenging cases.
@@
where $\hat{e}$ and $e^{*}$ represent two code snippets (LLM output and ground truth, respectively), $ed(\hat{e},e^{*})$ is computed using the classical dynamic programming algorithm (lcvenshtcin1966binary), and $l(\cdot)$ returns the length of a code snippet.
@@
GRPO is one of the most widely adopted methods in both industry and research within the field of reinforcement learning (pennino2025reasoning; li2025reinforcement). It is particularly effective for addressing the optimization problem in Eq. 1. A key advantage of GRPO is that it eliminates the need for a separate value function approximation; instead, it calculates advantages using the average reward across multiple sampled responses (rollouts) generated from the same input prompt. When using GRPO from an RL view, we can rewrite Eq. 1 to be
@@
where $\pi_{\theta}$, $\pi_{\theta_{old}}$, and $\pi_{ref}$ are the updating policy model, old policy model, and reference model, respectively. For each prompt $q$, GRPO samples a group of $G$ edits $\{e_{1},e_{2},\ldots,e_{G}\}$ from the old policy $\pi_{\theta_{old}}$. Here, $\epsilon$ and $\beta$ are hyperparameters that control the optimization behavior, and $\mathbb{D}_{\mathrm{KL}}$ denotes the KL divergence between the current training policy and a fixed reference policy (see (shao2024deepseekmath) for details).
@@
As demonstrated in our experiments (see Figure 7 for details), a substantial fraction of input prompts in real-world code editing scenarios yield skewed reward distributions. This skew arises from the inherent noise, complexity, and dynamism of practical environments—often manifested through outliers—which are unavoidable and unpredictable due to factors such as rich contextual dependencies, historical user behaviors, evolving intents, and other confounding variables. These outliers introduce noise into RL training (frauenknechtrollouts; hollenstein2022action) when using Eq. 4, since Eq. 4 employs an aggregation function that uniformly considers all reward values. Common approaches to mitigating the influence of outliers include using quantile-based statistics (e.g., Q1, median, Q3), which are more robust than the mean in the presence of outliers (john2015robustness; rousseeuw2011robust). However, as shown in Figure 7 and Figure 8, real reward distributions across different input prompts vary—being left-skewed, normal, or right-skewed—so a single quantile or mean cannot adapt effectively to diverse cases. To address this issue, we propose a group-adaptive advantage calculation approach to enhance existing GRPO and its variants (such as DAPO (yu2025dapo)); this novel RL method is termed Group Adaptive Policy Optimization (GAPO). The core idea is to identify the highest-density region of rewards, where most reward values concentrate, and the influence of outliers is minimal.
@@
Our GAPO method does not alter the objective of existing group-relative RL approaches; instead, it only modifies the advantage computation in Eq. 4. This design makes it simple to implement and plug-and-play with any RL framework, such as verl (sheng2025hybridflow), requiring only a few lines of code. Specifically, we redefine the advantage as
@@
To identify an outlier-free region that adapts to different reward distributions, we seek a sub-region with the highest density, which represents the majority of the data while ignoring outliers. This naturally leads us to a classical statistical problem: finding the highest-density interval (HDI) for a given probability mass (o2022smallest), which also identifies the narrowest interval. Formally, for a set of reward values, this can be expressed as
@@
For comprehensive evaluation, we consider nine large language models (LLMs) covering a diverse spectrum: Mistral-v0.3 (jiang2023mistral7b) (general-purpose, non-reasoning; 7B), Qwen2.5 (team2024qwen2) (general-purpose, non-reasoning; 3B, 7B), Qwen3 (yang2025qwen3) (general-purpose, reasoning; 4B, 8B), Qwen2.5-Coder (hui2024qwen2) (code-specialized, non-reasoning; 3B, 7B, 14B), and DeepSeek-Coder (guo2024deepseek) (code-specialized, non-reasoning; 6.7B). All these models are instruction-tuned versions downloaded from Hugging Face (huggingface).
@@
Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) has emerged as one of the most influential RL algorithms for LLM post-training (kumar2025llm; patil2025advancing; lai2025survey), widely adopted in both academia and industry. Recently, Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) (yu2025dapo) has gained attention as a promising extension and popular successor to GRPO. Our proposed GAPO method focuses on enhancing the group advantage computation—the core mechanism of GRPO. Accordingly, we adopt both the classical GRPO and the newly introduced DAPO as our primary baselines. We demonstrate here that the adaptive advantage computation, at the heart of GAPO, is a generalizable framework that can be seamlessly integrated into both GRPO and DAPO, leading to consistent performance improvements.
@@
We use the popular verl framework (sheng2025hybridflow) for RL training, which is widely adopted in industry due to its scalability. We follow most of the default settings for GRPO and DAPO in verl with adaptive modifications for our code edit tasks, such as an input prompt length of 4096, an output length of 1024, a rollout batch size of 512, a training batch size of 32, 8 rollouts per iteration, and a total of 10 epochs. Our GAPO method is straightforward to implement, requiring only a few lines of code by modifying the compute_grpo_outcome_advantage function in the GRPO and DAPO implementations (details are provided in our released code), demonstrating its high compatibility. We use the exact match ($em$) as the evaluation metric. All results are reported as the average over three trials.
@@
We investigate the reward distributions of rollouts for all input prompts in the training set, with the resulting statistics shown in Figure 7. The pie chart indicates that most input prompts (68.4%) induce rollouts with approximately normal distributions, while a substantial portion (19.9%) exhibits skewed distributions, roughly evenly split between left-skewed and right-skewed. This skew arises from the inherent noise, complexity, and dynamism of real-world environments—often manifesting as unavoidable and unpredictable outliers driven by complex context, user history, evolving intents, and other confounding factors. When computing advantage values as in Eq. 4 for these skewed distributions, outliers can have a negative impact, introducing noise (frauenknechtrollouts; hollenstein2022action). In contrast, our GAPO method employs an adaptive $Q$ value to locate the center of the dense region in the reward distribution, thereby mitigating the influence of outliers and reducing noise in the computation of advantage values.
@@
Furthermore, we visualize representative examples sampled from input prompts that exhibit four types of reward distributions in Figure 8, providing a closer look at the skewed cases. Noticeable gaps exist between the mean and median for the left-skewed, right-skewed, and approximately symmetric distributions, while outliers primarily appear in the left- and right-skewed cases. By further analyzing the relationship between the mean–median difference and the sign of the advantages computed in Eq. 4 and Eq. 5, we find that our GAPO method generates more negative rollouts than the original GRPO/DAPO in left-skewed distributions. This occurs because $Q$ is larger than the mean in such cases, causing most rewards to fall below $Q$ and thus produce negative advantages. This behavior is reasonable, as left-skewed distributions typically correspond to relatively easy problems with generally higher rewards, where more negative samples promote better generalization during training (mu2025dissecting; zhu2025surprising). Conversely, for right-skewed distributions (relatively hard problems), GAPO encourages more specialized learning, leading to improved accuracy on these challenging cases. These trends align with our post-training goals.
@@
To address the instability and limited exploration of GRPO, recent studies have proposed improvements from multiple perspectives. For instance, (wei2025redit) highlights GRPO’s gradient instability and mitigates it by dithering discrete reward signals with random noise. GMPO (zhao2025geometric) replaces the arithmetic mean with a geometric mean to reduce outlier sensitivity. Dr. GRPO (liu2025understanding) corrects training bias by jointly considering response length and question difficulty. GSPO (zheng2025group) mitigates high variance in Mixture-of-Experts models via sequence-level importance sampling. Other works (cui2025entropy; hao2025rethinking) stabilize training and alleviate entropy collapse by regulating entropy dynamics.
Beyond token-level optimization, another line of work enhances GRPO stability through refined advantage computation. OPO (hao2025policy) achieves stable policy optimization via a variance-minimizing bias term. BNPO (xia2025selection) adaptively normalizes rewards using a Beta distribution for low-variance gradient estimation. (wang2025kalman) employs Kalman filtering to estimate latent reward baselines and uncertainty. GRPO-MA (wang2025grpo) reduces variance by averaging multi-answer rewards, while (wu2025quantile) introduces a group-wise $K$-quantile baseline for entropy-safe reasoning.
@@
Code editing (li2023codeeditor; guo2024codeeditorbench; nam2025prompting) is more challenging than basic code refinement, as it depends not only on the user prompt but also on the coding context, cursor position, and historical edits. To tackle this challenge, Self-Edit (zhang2023self) fine-tunes a fault-aware neural editor in a generate-and-edit manner, improving code quality and accuracy. EDITLORD (li2025editlord), in contrast, avoids direct prompting or fine-tuning; it first uses an LLM to extract transformation rules from training data and then trains a second LLM to apply these rules to code pairs. IterPref (wu2025iterpref) leverages offline preference learning for iterative debugging, enabling context-aware code editing based on user feedback. LEDEX (jiang2024ledex) automatically collects refinement trajectories and enhances LLMs’ self-debugging ability via supervised fine-tuning and reinforcement learning. Finally, RLEF (gehring2024rlef) performs code editing with real-time execution feedback, optimizing refinement through end-to-end reinforcement learning.